# Learning Exactly Linearizable Deep Dynamics Models

### v1

@@ Appendix 1. Exactly Linearizable Condition (Su, 1982)
$\frac{dy}{dt}=\overline{f}(y)+\overline{g}(y)v.$ (36)
However, we consider a single-input system ($v\in\mathbb{R}$) for simplicity. Exact linearization implies that we can transform (36) into the following linear system by applying a coordinate transformation (4) and feedback (5)
$\frac{dx}{dt}=\begin{pmatrix}0&1&0&\cdots&0\\ 0&0&1&\cdots&0\\ \vdots&\vdots&&\ddots&0\\ 0&0&0&\cdots&1\\ 0&0&0&\cdots&0\end{pmatrix}x+\begin{pmatrix}0\\ 0\\ \vdots\\ 0\\ 1\end{pmatrix}u.$ (47)
@@
$\left[\overline{f},\overline{g}\right]$ $:=\frac{\partial\overline{g}}{\partial y}\overline{f}(y)-\frac{\partial\overline{f}}{\partial y}\overline{g}(y).$ (49)
@@
$\mathbf{ad}^{0}_{\overline{f}}\overline{g}$ $=\overline{g}(y),$ (51)
$\mathcal{L}_{\mathbf{ad}_{\overline{f}}\overline{g}}\phi(y)$ $=\mathcal{L}_{\overline{f}}\mathcal{L}_{\overline{g}}\phi(y)-\mathcal{L}_{\overline{g}}\mathcal{L}_{\overline{f}}\phi(y),$ (52)
where $\phi(y)$ denotes the arbitrary function $y$.
@@
    $\left\{\mathbf{ad}^{0}_{\overline{f}}\overline{g},\mathbf{ad}^{1}_{\overline{f}}\overline{g},\cdots,\mathbf{ad}^{n-1}_{\overline{f}}\overline{g}\right\}(y)$ is linearly independent for any $y$.
@@
    $\left\{\mathbf{ad}^{0}_{\overline{f}}\overline{g},\mathbf{ad}^{1}_{\overline{f}}\overline{g},\cdots,\mathbf{ad}^{n-2}_{\overline{f}}\overline{g}\right\}(y)$ is involutive.
@@ Appendix 2. Nonlinear Optimal Control for Exactly Linearizable Input-Affine Dynamics
We proposed a control design method that fully utilizes simple linear optimal control theory. In this section, we show how to utilize nonlinear optimal control theory based on the modeling framework in the present study, assuming that the system to be controlled is an input-affine nonlinear system. Let us consider the following input-affine nonlinear system, as in (4) and (5). We can construct $\Psi_{1},\ \Psi_{2},\ A,\ B,\ c$ as
$v(t)$ $=\Psi_{1}(y(t))+\Psi_{2}(y(t))u(t),$ (53)
$\dot{x}(t)$ $=Ax(t)+Bu(t)+c,$ (54)
$y(t)$ $=\Phi^{-1}(x(t)),$ (55)
using the data for $\{v,y,\dot{y}\}$, where $\Phi$ is bijective and $\Psi_{2}(y)$ is nonsingular for any $y$. Without the loss of generality, we assume $\Phi(0)=0$ and $c=0$. The obtained dynamics are input affines in the form of (3). Next, we define the control Lyapunov function (CLF), which is essential for explaining the Sontag-type stabilizing control law.
@@
$\inf_{v\in D}\left\{\mathcal{L}_{\bar{f}}V(y)+\mathcal{L}_{\bar{g}}V(y)v\right\}<-\gamma(V(y)).$ (56)
@@
${V}(y)=\Phi(y)^{\mathrm{T}}P\Phi(y)$ (57)
with the solution $P$ to (20). Once a CLF is found, we utilize several tools from the nonlinear control theory.
@@
Let a positive definite function $V:Y\rightarrow\mathbb{R}$ be a CLF for system (3); then, the following Sontag-type control law $u=\alpha_{d}(y)$ asymptotically stabilizes the origin of the system (3).
$u$ $=\alpha_{d}(y)$ (58)
$\alpha_{d}(y)$ :=$:=\begin{cases}-\frac{1}{r_{d}(y)}(\mathcal{L}_{g}V)^{\mathrm{T}},\ \ &(\mathcal{L}_{g}V\neq 0)\\ 0,&(\mathcal{L}_{g}V=0)\end{cases}$ (59)
where the bounded function $r_{d}(y)>0$ is given by
$r_{d}(y):=\frac{\mathcal{L}_{g}V(\mathcal{L}_{g}V)^{\mathrm{T}}}{\mathcal{L}_{f}V+\sqrt{\mathcal{L}_{f}V^{2}+(\mathcal{L}_{g}V(\mathcal{L}_{g}V)^{\mathrm{T}})^{2}}}.$ (60)
@@ Appendix 3. Equilibrium Condition of Proposed Control System
$\frac{\mathrm{d}}{\mathrm{d}t}\begin{pmatrix}x\\ u\\ \end{pmatrix}=\begin{pmatrix}f(x)+g(x)u\\ \lambda\\ \end{pmatrix},$ (65)
$\begin{cases}\lambda^{\ast}=&\mathop{\mathrm{arg~{}min}}\limits_{\lambda}\frac{\mathrm{d}}{\mathrm{d}t}\|u-k(x)\|^{2}+a\|\lambda\|^{2}\\ \mathrm{s.t.}&\frac{\partial h}{\partial x}(f(x)+g(x)u)+\frac{\partial h}{\partial u}\lambda\leq\alpha(-h(x,u))\end{cases},$ (66)
where $a>0$ and $h(x,u)$ are convex with respect to $u$, for any fixed $x$. The convexity originated from the structure of the EL model. We can prove the optimality of the equilibrium as
@@
If an equilibrium point $(x,u)$ exists in the system shown in (65) and (66), then the points satisfy $u=\mathop{\mathrm{arg~{}min}}\limits_{u}\|u-k(x)\|^{2}\ \mathrm{s.t.}\ h(x,u)\leq 0$.
@@
As the problem in (66) is strongly convex with $a>0$, the necessary and sufficient condition for first-order optimality can be written as
@@
where $F(\lambda):=\frac{\mathrm{d}}{\mathrm{d}t}\|u-k(x)\|^{2}+a\|\lambda\|^{2},\ G(\lambda):=\frac{\partial h}{\partial x}(f(x)+g(x)u)+\frac{\partial h}{\partial u}\lambda-\alpha(-h(x,u))$, and $\mu$ represents a Lagrange multiplier. Considering the equilibrium condition of (65), i.e., $f(x)+g(x)u=0,\ \lambda=0$, the above optimality condition reduces to
$2(u-k(x))+{\left(\frac{\partial h}{\partial u}\right)}^{\mathrm{T}}\mu=0,$ (69)
$\mu\geq 0,\ h(x,u)\leq 0,\ \mu\circ h(x,u)=0.$ (70)
$h(x,u)$ is convex to $u$, and therefore, the above condition coincides with the necessary and sufficient condition of the first-order optimality of the optimization problem
$\mathop{\mathrm{minimize}}\limits_{u}\|u-k(x)\|^{2}\ \mathrm{s.t.}\ h(x,u)\leq 0.$ (71)
v>

### v2

@@ 5 Conclusion
Notes on contributors
Ryuta Moriyasu received his B.E. and M.E. degrees from Osaka University in 2011 and 2013, and his Ph.D. degree from Kyoto University in 2024. Since 2013, he has been with Toyota Central R&D Labs., Inc., Aichi, Japan. His research interests include machine learning, control theory, optimization theory, and their applications in the automotive field. Recently, he has been working on theoretical reliability assurance in control design using machine learning models, with the aim of practical application to engine control system. He received Best Presentation Award of the Multi-Symposium on Control System from SICE Control Division, Excellent Presentation Award from RSJ, The Asahara Science Award from JSAE, and so on. He is a Member of SICE and JSAE.
Masayuki Kusunoki received his B.E. and M.E. degrees from Shizuoka University in 2013 and 2015. Since 2015, he has been with Toyota Industries Corporation, Aichi, Japan. He has been engaging in control development of diesel engines for passenger cars.
Kenji Kashima received his Doctoral degree in Informatics from Kyoto University in 2005. He was with Tokyo Institute of Technology, Universität Stuttgart, Osaka University, before he joined Kyoto University in 2013, where he is currently an Associate Professor. His research interests include control and learning theory for complex dynamical systems, and their applications. He received Humboldt Research Fellowship (Germany), IEEE CSS Roberto Tempo Best CDC Paper Award, Pioneer Award of SICE Control Division, and so on. He has served as an Associate Editor of IEEE Transactions of Automatic Control, IEEE CONTROL SYSTEMS LETTERS, and IEEE CSS Conference Editorial Board and a steering committee member of MTNS. He is a Senior Member of IEEE and Member of SICE, ISCIE and IEICE.
@@ Appendix 1. Exactly Linearizable Condition (Su, 1982)
$\frac{dx}{dt}=\overline{f}(x)+\overline{g}(x)u.$ (37)
However, we consider a single-input system ($u\in\mathbb{R}$) for simplicity. Exact linearization implies that we can transform (37) into the following linear system by applying a coordinate transformation (4) and feedback (5)
$\frac{d\xi}{dt}=\begin{pmatrix}0&1&0&\cdots&0\\ 0&0&1&\cdots&0\\ \vdots&\vdots&&\ddots&0\\ 0&0&0&\cdots&1\\ 0&0&0&\cdots&0\end{pmatrix}\xi+\begin{pmatrix}0\\ 0\\ \vdots\\ 0\\ 1\end{pmatrix}v.$ (38)
@@
$\left[\overline{f},\overline{g}\right]$ $:=\frac{\partial\overline{g}}{\partial x}\overline{f}(x)-\frac{\partial\overline{f}}{\partial x}\overline{g}(x).$ (40)
@@
$\mathbf{ad}^{0}_{\overline{f}}\overline{g}$ $=\overline{g}(x),$ (42)
$\mathcal{L}_{\mathbf{ad}_{\overline{f}}\overline{g}}\phi(x)$ $=\mathcal{L}_{\overline{f}}\mathcal{L}_{\overline{g}}\phi(x)-\mathcal{L}_{\overline{g}}\mathcal{L}_{\overline{f}}\phi(x),$ (43)
where $\phi(x)$ denotes the arbitrary function $x$.
@@
    $\left\{\mathbf{ad}^{0}_{\overline{f}}\overline{g},\mathbf{ad}^{1}_{\overline{f}}\overline{g},\cdots,\mathbf{ad}^{n-1}_{\overline{f}}\overline{g}\right\}(x)$ is linearly independent for any $x$.
@@
    $\left\{\mathbf{ad}^{0}_{\overline{f}}\overline{g},\mathbf{ad}^{1}_{\overline{f}}\overline{g},\cdots,\mathbf{ad}^{n-2}_{\overline{f}}\overline{g}\right\}(x)$ is involutive.
@@ Appendix 2. Nonlinear Optimal Control for Exactly Linearizable Input-Affine Dynamics
We proposed a control design method that fully utilizes simple linear optimal control theory. In this section, we show how to utilize nonlinear optimal control theory based on the modeling framework in the present study, assuming that the system to be controlled is an input-affine nonlinear system.
Let us consider the following input-affine nonlinear system, as in (4) and (5). We can construct $\Psi_{1},\ \Psi_{2},\ A,\ B,\ c$ as
$u(t)$ $=\Psi_{1}(x(t))+\Psi_{2}(x(t))v(t),$ (44)
$\dot{\xi}(t)$ $=A\xi(t)+Bv(t)+c,$ (45)
$x(t)$ $=\Phi^{-1}(\xi(t)),$ (46)
using the data for $\{u,x,\dot{x}\}$, where $\Phi$ is bijective and $\Psi_{2}(x)$ is nonsingular for any $x$. Without the loss of generality, we assume $\Phi(0)=0$ and $c=0$. The obtained dynamics are input-affine in the form of (3). Next, we define the control Lyapunov function (CLF), which is essential for explaining the Sontag-type stabilizing control law.
@@
$\inf_{v\in D}\left\{\mathcal{L}_{\bar{f}}V(x)+\mathcal{L}_{\bar{g}}V(x)u\right\}<-\gamma(V(x)).$ (47)
@@
${V}(x)=\Phi(x)^{\mathrm{T}}P\Phi(x)$ (48)
with the solution $P$ to (27). Once a CLF is found, we can utilize several tools from the nonlinear control theory.
@@
Let a positive definite function $V:Y\rightarrow\mathbb{R}$ be a CLF for system (3); then, the following Sontag-type control law $v=\alpha_{d}(x)$ asymptotically stabilizes the origin of the system (3).
$v$ $=\alpha_{d}(x)$ (49)
$\alpha_{d}(x)$ :=$:=\begin{cases}-\frac{1}{r_{d}(x)}(\mathcal{L}_{g}V)^{\mathrm{T}},\ \ &(\mathcal{L}_{g}V\neq 0)\\ 0,&(\mathcal{L}_{g}V=0)\end{cases}$ (50)
where the bounded function $r_{d}(x)>0$ is given by
$r_{d}(x):=\frac{\mathcal{L}_{g}V(\mathcal{L}_{g}V)^{\mathrm{T}}}{\mathcal{L}_{f}V+\sqrt{\mathcal{L}_{f}V^{2}+(\mathcal{L}_{g}V(\mathcal{L}_{g}V)^{\mathrm{T}})^{2}}}.$ (51)
@@ Appendix 3. Equilibrium Condition of Proposed Control System
$\frac{\mathrm{d}}{\mathrm{d}t}\begin{pmatrix}\xi\\ v\\ \end{pmatrix}=\begin{pmatrix}f(\xi)+g(\xi)v\\ \lambda\\ \end{pmatrix},$ (52)
$\begin{cases}\lambda^{\ast}=&\mathop{\mathrm{arg~{}min}}\limits_{\lambda}\frac{\mathrm{d}}{\mathrm{d}t}\|v-k(\xi)\|^{2}+\beta\|\lambda\|^{2}\\ \mathrm{s.t.}&\frac{\partial h}{\partial\xi}(f(\xi)+g(\xi)v)+\frac{\partial h}{\partial v}\lambda\leq\alpha(-h(\xi,v))\end{cases},$ (53)
where $\beta>0$ and $h(\xi,v)$ are convex with respect to $v$, for any fixed $\xi$. The convexity originated from the structure of the EL model. We can prove the optimality of the equilibrium as
@@
If an equilibrium point $(\xi,v)$ exists in the system shown in (52) and (53), then the points satisfy $v=\mathop{\mathrm{arg~{}min}}\limits_{v}\|v-k(\xi)\|^{2}\ \mathrm{s.t.}\ h(\xi,v)\leq 0$.
@@
As the problem in (53) is strongly convex with $\beta>0$, the necessary and sufficient condition for first-order optimality can be written as
@@
where $F(\lambda):=\frac{\mathrm{d}}{\mathrm{d}t}\|v-k(\xi)\|^{2}+\beta\|\lambda\|^{2},\ G(\lambda):=\frac{\partial h}{\partial\xi}(f(\xi)+g(\xi)v)+\frac{\partial h}{\partial v}\lambda-\alpha(-h(\xi,v))$, and $\mu$ represents a Lagrange multiplier. Considering the equilibrium condition of (52), i.e., $f(\xi)+g(\xi)v=0,\ \lambda=0$, the above optimality condition reduces to
$2(v-k(\xi))+{\left(\frac{\partial h}{\partial v}\right)}^{\mathrm{T}}\mu=0,$ (56)
$\mu\geq 0,\ h(\xi,v)\leq 0,\ \mu\circ h(\xi,v)=0.$ (57)
$h(\xi,v)$ is convex to $v$, and therefore, the above condition coincides with the necessary and sufficient condition of the first-order optimality of the optimization problem
$\mathop{\mathrm{minimize}}\limits_{v}\|v-k(\xi)\|^{2}\ \mathrm{s.t.}\ h(\xi,v)\leq 0.$ (58)
Word Count
This manuscript includes 4424 words.